CARA Telecon Minutes Date: March 5, 2001 Time: 1:00 PM Attendance: Boston University: Cornell: CMU: J. Peterson Northwestern RIT: SAO: A. Lane, T. Stark South Pole: C.Martin, C. Kaminski University of Chicago: J. Carlstom, A.Harper, S. Meyer, N. Odalen R. Pernic, R. Landsberg UCB: UCSB: J. Ruhl Approval of previous minutes: It was requested that CARA members review the previous minutes and send changes/corrections and approval via email to John Carlstrom. Agenda Items Projects News: DASI - everything fine - Compressor room had cooled due to weather conditions at Pole. Now back closer to normal. AST/RO - 230 & 460 GHz receiver is working. Both are observing, full beam procession on the sky. - Weather good for day or two. - FTS broke - has too much friction. Repairs should be minor. FTS has been cold and running continuously since Pole closed. This is different than reported last week -- urged NOT to cool with liquid helium until weather is better and liquid helium situation is understood and optimal use plan developed. J. Carlstrom suggested that the use of the FTS be postponed until an overall CARA plan for the winter liquid helium usage is developed -- since the FTS uses a great deal of helium. For the time being, with the weather not good, we should perhaps not run the system. ACBAR - Taking lots of data. No problems with freeze cycle. - Loss of helium in transfers is much lower after consultation. - Now in process of different observing strategy. Going over some previous clusters. Still observing clusters with sun up. - Accumulating data at Berkeley, data transmission going well. - Slow link out of White Sands have requested CARA to have larger bandwidth from White Sands. Limiting factor is the phone line out of White Sands *** this is action item for Bob Loewenstein at upcoming projects meeting. *** News from the Pole: C. Kaminski reported that the Marisat is having problems this year. Every time satellite is about to make pass at Pole someone has to go out to the berm to press some buttons. . . ACTION ITEM - R. Loewenstein to discuss communication issues during projects meeting in March. Helium at Pole: J. Ruhl spoke with Wessington and Raytheon and came up with plan. Plan is to go to normal Wessington configuration of neck. Raytheon sent out email to A. Baker and C. Kaminski to do this to Wesintgon number one or two and see how things go. Latest numbers were sent by A. Barker but had not been reviewed by the CARA group. C. Kaminski reported that the change had not taken place on the Wessington's that the plan was to modify today. Delay was caused by problems at MAPO with the heat. MAPO dropped to 40F. Modification of the pressure done early on the Wessington was only a temporary gain. The other dewar has had its pressure lowered. Projects Meeting: - Begin with science talks - results from the instruments. - Look over the agenda to see if anything is missing and contact J. Carlstrom with information. - R. Pernic is working on the SIPS. They are due in 3 weeks. Be prepared to give R. Pernic information for your project. - J. Carlstrom and S. Meyer will be talking with NSF-OPP next Wednesday and will report at the projects meeting. If you want them to speak on a particular issue contact J. Carlstrom. - Please respond to N. Odalen regarding email that will be sent on the meeting logistics. REU: We have only heard from one institution on their need for a student for the summer. A. Lane reported that they have positions available at SAO. The group needs to respond to R. Landsberg as soon as possible as students are being offered other opportunities.